A ball is dropped from a height h on a floor of coefficient of restitution e. The total distance covered by ball just after second hit . Please tell about distance covered after n ^th hit

Grade:12th pass

2 Answers

Arun
5 years ago
Dear student
 
As we know
 
e = sqrt (h1/h)
 
Hence
 
e² = h1/ h
 
h1 = e² *h
 
 
Before 2nd hit
 
Distance travelled = h + 2 h1
 = h + 2 e²h
 = h( 1 +2 e²)
Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your problem.
 
As we know
e = (h1/h)
Hence
e² = h1/ h
h1 = e² *h
 
Before 2nd hit
Distance travelled = h + 2 h1
 = h + 2 e²h
 = h( 1 +2 e²)
 
Thanks and regards,
Kushagra